# Local Dissipativity Analysis of Nonlinear Systems

### v1

@@ 4.2 Analysis via Optimization
The conditions for a CPA or combined quadratic and CPA function to adhere to the dissipativity LMI (5) developed in Theorems 12 and 13, respectively, provide the foundation to formulate optimization problems for local dissipativity analysis. Problem 4.1 and Problem 4.2 describe a general optimization problem to determine the dissipativity of a system while simultaneously synthesizing a storage function. Problem 4.1 considers systems with no linear control affine terms, while Problem 4.2 considers those with linear control affine terms. In both problems, the generic cost function $J(\alpha,\beta)$ is used, where $\alpha$ and $\beta$ can be determined based of the desired QSR characterization being analyzed. These options are given in Table 1.
@@
Consider the mapping $\mathcal{G}_{\Omega}(x_{0}):\mathcal{U}_{\Omega}\rightarrow\mathcal{L}_{2e}^{p}$ defined by $\mathbf{y}=\mathcal{G}_{\Omega}(x_{0})\mathbf{u}$ and (1). Let $\mathbf{B}=\mathbf{0}$ and $\mathbf{D}=\mathbf{0}$. Define the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ over $\Omega\in\mathfrak{R}^{n},$ the CPA function $V=\{V_{\mathbf{x}}\}_{\mathbf{x}\in\mathbb{E}_{\mathcal{T}}}\in\mathbb{R}^{n}$, and the variables $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R}$, $\alpha\in\mathbb{R},$ and $\beta\in\mathbb{R}.$ Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal matrices. The convex optimization problem is defined as
@@
$\min_{\alpha,\beta,V,\mathbf{L}}J(\alpha,\beta)$
@@
IO Characterization QSR Design Variables Objective (0.1)
$\mathcal{L}_{2}$-gain $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\mathbf{0}$, $\mathbf{R}=\gamma^{2}\mathbf{I}$ $\alpha=\gamma^{2}$, $\beta=0$ $J(\alpha,\beta)=\alpha$ $\alpha>0$
Input Strictly Passive $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=-\nu\mathbf{I}$ $\alpha=\nu$, $\beta=0$ $J(\alpha,\beta)=-\alpha$ $\alpha>0$
Output Strictly Passive $\mathbf{Q}=-\rho\mathbf{I}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=\mathbf{0}$ $\alpha=\rho$, $\beta=0$ $J(\alpha,\beta)=-\alpha$ $\alpha>0$
Nondegenerate conic [joshi2002design, bridgeman2014conic] $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\frac{a+b}{2}\mathbf{I},$ $\mathbf{R}=-ab\mathbf{I}$ $\alpha=\frac{a+b}{2},$ $\beta=ab$ $J(\alpha,\beta)=\alpha^{2}-2\beta$ $\beta<0$
Degenerate interior conic [bridgeman2014conic] $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\mathbf{I}$, $\mathbf{R}=-d\mathbf{I}$ $\alpha=d$ $J(\alpha,\beta)=-\alpha$ $\alpha<0$
QSR dissipative $\mathbf{Q}\preceq 0,$ any $\mathbf{S}$, $\mathbf{R}\succeq 0$ $\alpha=0$, $\beta=0$ $J(\alpha,\beta)=0$
@@ 5 Solution guarantees
Finally, we establish that we are able to determine a CPA or quadratic storage function for a system that is locally dissipative. It is important to note that the QSR dissipativity holds with some amount of strictness. Theorems 12 and 13 uses error bounds to ensure the validity of a QSR dissipativity characterization. Therefore, the solution found by Problems 1 and 2 will always be conservative at some level.
@@
Consider Problem 4.1. Let $\mathcal{G}_{\Omega}$ be locally QSR dissipative on $\Omega$ with $\mathbf{Q}\preceq-\frac{1}{2}(\Pi_{2}+\Pi_{3})^{-1}$ and $\hat{\mathbf{R}}\preceq-\frac{1}{2}(\Pi_{1}^{-1}+\eta\mathbf{I})$, where $\eta\geq\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\hat{J}_{i,j}^{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}$ and $j\in\mathbb{Z}_{0}^{n}.$ Let there exist a storage function $\phi\in\mathcal{C}^{2}$ for which (5) holds in $\Omega,$ where $\sup_{\mathbf{x}\in\Omega}\left\lVert\nabla\phi\right\rVert_{2}\leq C$ and $\max_{i,j\in\mathbb{Z}_{1}^{n},\mathbf{x}\in\Omega}\left\lvert\frac{\partial\phi}{\partial\mathbf{x}^{(i)}\partial\mathbf{x}^{(j)}}\Big|_{\mathbf{x}}\right\rvert\leq A$ for some $C,A<\infty.$ Then, Problem 4.1 will find a valid CPA storage function over $\Omega$ for (1) using a finite number of simplices.
@@
Let $\bar{t}$ be the maximum length of any simplex in $\mathcal{T}.$ On each simplex $\sigma_{i}$, define
$\bar{\phi}_{i}=\begin{bmatrix}\phi(\mathbf{x}_{i,1})-\phi(\mathbf{x}_{i,0})\\ \vdots\\ \phi(\mathbf{x}_{i,n})-\phi(\mathbf{x}_{i,0})\end{bmatrix}$,
and define $\mathbf{X}_{i}$ as in Lemma 7.
Let each vertex of the function $V$ be defined by the storage function, $\phi$, i.e $V(x_{i,j})=\phi(x_{i,j})$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ $j\in\mathbb{Z}_{0}^{n}.$ By Lemma 7, the gradient of the CPA function $\nabla V_{i}$ on each simplex is $\mathbf{X}_{i}^{-1}\bar{\phi}.$
@@
$\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}=$ $\frac{1}{2}\left\lVert\begin{bmatrix}\overline{\mathbf{H}}_{\phi}(x_{i,1})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\\ \vdots\\ \overline{\mathbf{H}}_{\phi}(x_{i,n})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\end{bmatrix}\right\rVert_{2}$
$=$ $\frac{1}{2}\sqrt{\sum_{j=1}^{n}\left\lvert\overline{\mathbf{H}}_{\phi}(x_{i,j})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\right\rvert^{2}}.$
By lemma with $\mathbf{H}$ bound, this is bounded above by $An^{\frac{3}{2}}\bar{t}^{2}.$ Thus, $\left\lVert\mathbf{X}_{i}^{-1}\bar{\phi}-\nabla\phi(x)\right\rVert_{2}$
@@
Consider Constraint (14d).For any element of the matrix that contains $\nabla V$, add and subtract $\nabla\phi,$ e.g. $\nabla V^{\top}f(x_{i,j})=(\nabla V-\nabla\phi+\nabla\phi)^{\top}f(x_{i,j}).$ For each simplex, $\nabla V=\mathbf{X}_{i}^{-1}\bar{\phi}.$ Then, by the Schur complement [boyd1994linear], Constraint (14d) is equivalent to the inequality,
$\Lambda+(\mathbf{X}^{-1}\bar{\phi}-\nabla\phi)^{\top}f(x_{i,j})+\frac{1}{2}(\mathbf{X}^{-1}\bar{\phi}-\nabla\phi)^{\top}G(x_{i,j})\tilde{\mathbf{R}}^{-1}$
@@
$\tilde{\mathbf{R}}^{-1}(\frac{1}{2}G(x_{i,j})^{\top}\nabla\phi-\mathbf{S}^{\top}h(x_{i,j})-J(x_{i,j})^{\top}\tilde{\mathbf{Q}}h(x_{i,j})),$
which is equivalent to the (5) evaluated at $x_{i,j}$ using the true storage function, $\phi.$ Further, $\tilde{\mathbf{R}}=\hat{\mathbf{R}}+\frac{1}{2}(\Pi_{1}^{-1}+\hat{s}\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}),$ and $\tilde{\mathbf{Q}}=(\mathbf{Q}^{-1}+\frac{1}{2}\Pi_{2}^{-1}+\frac{1}{2}\Pi_{3}^{-1})^{-1},$ and $\hat{f}_{i,j},$ $\hat{h}_{i,j}$, $\hat{G}_{i,j}$, and $\hat{s}$ are defined by Theorem 12.
@@
$\Lambda{+}\Bigl(\left\lVert f(x_{i,j})\right\rVert_{2}{+}\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert G(x_{i,j}\right\rVert_{2}$
$+\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert h(x_{i,j})\right\rVert_{2}(\left\lVert\mathbf{S}\right\rVert+\left\lVert\tilde{\mathbf{Q}}\right\rVert_{2}\left\lVert J(x_{i,j}\right\rVert_{2})\Bigr)\Bigl(\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}$
$\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}\Bigr)+\hat{f}_{i,j}+\frac{1}{2\underline{\pi}_{1}}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j}+\frac{1}{4\underline{\pi}_{1}}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}$
$+\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}\hat{h}_{i,j}.$

### v2

@@ 4.2 Analysis via Optimization
The conditions for a CPA or combined quadratic and CPA function to adhere to (5) developed in Theorems 12 and 13, respectively, provide the foundation to formulate optimization problems for local dissipativity analysis. Problem 4.1 and Problem 4.2 describe a general optimization problem to determine the dissipativity of a system while simultaneously synthesizing a storage function. Problem 4.1 considers systems with no linear control affine terms, while Problem 4.2 considers those with linear control affine terms. In both problems, the generic cost function $J(\alpha,\beta)$ is used, where $\alpha$ and $\beta$ can be determined based of the desired QSR characterization being analyzed. Common options are given in Table 1.
@@
Consider the mapping $\mathcal{G}_{\Omega}(x_{0}):\mathcal{U}_{\Omega}\rightarrow\mathcal{L}_{2e}^{p}$ defined by $\mathbf{y}=\mathcal{G}_{\Omega}(x_{0})\mathbf{u}$ and (1). Let $\mathbf{B}=\mathbf{0}$ and $\mathbf{D}=\mathbf{0}$. Define the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ over $\Omega\in\mathfrak{R}^{n},$ the CPA function $V=\{V_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}\in\mathbb{R}^{n}$, and the variables $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R}$, $\alpha\in\mathbb{R},$ and $\beta\in\mathbb{R}.$ Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal matrices. The convex optimization problem is defined as
@@
$\min_{\alpha,\beta,V,\mathbf{L},\Pi}J(\alpha,\beta)$
@@
IO Characterization QSR Design Variables Objective
$\mathcal{L}_{2}$-gain $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\mathbf{0}$, $\mathbf{R}=\gamma^{2}\mathbf{I}$ $\alpha=\gamma^{2}$, $\beta=0$, $\alpha>0$ $J(\alpha,\beta)=\alpha$
Input Strictly Passive $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=-\nu\mathbf{I}$ $\alpha=\nu$, $\beta=0$, $\alpha>0$ $J(\alpha,\beta)=-\alpha$
Output Strictly Passive $\mathbf{Q}=-\rho\mathbf{I}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=\mathbf{0}$ $\alpha=\rho$, $\beta=0$, $\alpha>0$ $J(\alpha,\beta)=-\alpha$
Nondegenerate conic [joshi2002design, bridgeman2014conic] $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\frac{a+b}{2}\mathbf{I},$ $\mathbf{R}=-ab\mathbf{I}$ $\alpha=\frac{a+b}{2},$ $\beta=ab$, $\beta<0$ $J(\alpha,\beta)=\alpha^{2}-2\beta$
Degenerate interior conic [bridgeman2014conic] $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\mathbf{I}$, $\mathbf{R}=-d\mathbf{I}$ $\alpha=d$, $\alpha<0$ $J(\alpha,\beta)=-\alpha$
QSR dissipative (verification) $\mathbf{Q}\preceq 0,$ any $\mathbf{S}$, $\mathbf{R}\succeq 0$ $\alpha=0$, $\beta=0$ $J(\alpha,\beta)=0$
QSR dissipative (Selection) $\mathbf{Q}\preceq 0,$ any $\mathbf{S}$, $\mathbf{R}\succeq 0$ $\mathbf{Q},\ \mathbf{S},\ \mathbf{R}$ $J(\mathbf{Q},\mathbf{S},\mathbf{R})$ ==any convex function of its variables
@@ 5 Solution guarantees
Finally, we establish when a CPA or quadratic storage function is assured to be found for a locally dissipative system, as well as establishing when a combined quadratic and CPA storage function can be found. It is important to note that the QSR dissipativity holds with some amount of strictness. Theorems 12 and 13 use error bounds to ensure the validity of a QSR dissipativity characterization. Therefore, the solution found by Problems 1 and 2 will always be conservative at some level.
@@
Consider Problem 4.1. Let $\mathcal{G}_{\Omega}$ be locally QSR dissipative on a triangulable set, $\Omega$, with $\mathbf{Q}{\preceq}{-}\frac{1}{2}(\Pi_{2}{+}\Pi_{3})^{-1}$ and $\hat{\mathbf{R}}{\preceq}{-}\frac{1}{2}(\Pi_{1}^{-1}{+}\eta\mathbf{I})$, where $\eta\geq\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\hat{J}_{i,j}^{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}$ and $j\in\mathbb{Z}_{0}^{n}.$ Let there exist a storage function $\phi\in\mathcal{C}^{2}$ for which (5) holds in $\Omega,$ where $\sup_{x\in\Omega}\left\lVert\nabla\phi\right\rVert_{2}\leq C$ and $\max_{i,j\in\mathbb{Z}_{1}^{n},x\in\Omega}\left\lvert\frac{\partial\phi}{\partial x^{(i)}\partial x^{(j)}}\Big|_{x}\right\rvert\leq A$ for some $C,A<\infty.$ Then, Problem 4.1 will find a valid CPA storage function over $\Omega$ for (1) using a finite number of simplices.
This proof parallels [giesl2012construction, Theorem 3.3], which determines when a CPA Lyapunov function can be found for an exponentially stable system, by bounded the error between the true Lyapunov function and the CPA Lyapunov function. Here, we instead consider the QSR dissipativity condition (5) rather than Lyapunov stability conditions – bounding the error between the system’s true storage function and the CPA storage function. This requires consideration of a matrix inequality rather than a scalar inequality.
@@
For some triangulation, $\mathcal{T}$, of $\Omega$, let $\bar{t}$ be the maximum length of any simplex in $\mathcal{T}.$ On each simplex $\sigma_{i}$, define
$\bar{\phi}_{i}=\begin{bmatrix}\phi(x_{i,1})-\phi(x_{i,0})\\ \vdots\\ \phi(x_{i,n})-\phi(x_{i,0})\end{bmatrix}$,
and define $\mathbf{X}_{i}$ as in Lemma 7. Define also a candidate CPA storage function over $\mathcal{T}$ as $V(x_{i,j})=\phi(x_{i,j})$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ $j\in\mathbb{Z}_{0}^{n}.$ By Lemma 7, the gradient of the CPA function $\nabla V_{i}$ on each simplex is $\mathbf{X}_{i}^{-1}\bar{\phi}.$
@@
$\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}=$ $\frac{1}{2}\left\lVert\begin{bmatrix}\overline{\mathbf{H}}_{\phi}(x_{i,1},x)-\overline{\mathbf{H}}_{\phi}(x_{i,0},x)\\ \vdots\\ \overline{\mathbf{H}}_{\phi}(x_{i,n},x)-\overline{\mathbf{H}}_{\phi}(x_{i,0},x)\end{bmatrix}\right\rVert_{2}$
$=$ $\frac{1}{2}\sqrt{\sum_{j=1}^{n}\left\lvert\overline{\mathbf{H}}_{\phi}(x_{i,j},x)-\overline{\mathbf{H}}_{\phi}(x_{i,0}),x\right\rvert^{2}}.$
By the triangle inequality and the properties of matrix operator norms, this is bounded above by
$\frac{1}{2}\sqrt{\sum_{j=1}^{n}(\left\lVert\mathbf{H}_{\phi}(z_{i,1}\!)\right\rVert_{2}\!\left\lVert x_{i,1}{-}x\right\rVert_{2}^{2}\!{+}\!\left\lVert\mathbf{H}_{\phi}(z_{i,0}\!)\right\rVert_{2}\!\left\lVert x_{i,0}{-}x\right\rVert_{2}^{2})^{2}}$
By [giesl2012construction, Lemma 2.3], this is bounded above by $An^{\frac{3}{2}}\bar{t}^{2}.$ Thus, $\left\lVert\mathbf{X}_{i}^{-1}\bar{\phi}-\nabla\phi(x)\right\rVert_{2}$
@@
Consider the matrix $\mathbf{M}$ in Constraint (14d). For any element of $\mathbf{M}$ that contains $\nabla V$, add and subtract $\nabla\phi,$ e.g. $\nabla V^{\top}f(x_{i,j})=(\nabla V-\nabla\phi+\nabla\phi)^{\top}f(x_{i,j}).$ For each simplex, $\nabla V=\mathbf{X}_{i}^{-1}\bar{\phi}.$ Then, by the Schur complement [boyd1994linear], Constraint (14d) is equivalent to the inequality,
$\Lambda+(\mathbf{X}^{-1}\bar{\phi}{-}\nabla\phi)^{\top}f(x_{i,j})$
$+\frac{1}{2}(\mathbf{X}^{-1}\bar{\phi}{-}\nabla\phi)^{\top}G(x_{i,j})\tilde{\mathbf{R}}^{-1}$
@@
$\tilde{\mathbf{R}}^{-1}(\frac{1}{2}G(x_{i,j})^{\top}\nabla\phi{-}\mathbf{S}^{\top}\!h(x_{i,j}){-}J(x_{i,j})\!^{\top}\!\tilde{\mathbf{Q}}h(x_{i,j})).$
Evaluating (5) at $x_{i,j}$ using the asserted storage function, $\phi,$ equals $\Lambda$. Further, $\tilde{\mathbf{R}}=\hat{\mathbf{R}}+\frac{1}{2}(\Pi_{1}^{-1}+\hat{s}\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}),$ and $\tilde{\mathbf{Q}}=(\mathbf{Q}^{-1}+\frac{1}{2}\Pi_{2}^{-1}+\frac{1}{2}\Pi_{3}^{-1})^{-1},$ and $\hat{f}_{i,j},$ $\hat{h}_{i,j}$, $\hat{G}_{i,j}$, and $\hat{s}$ are defined by Theorem 12.
@@
$\Lambda{+}\Bigl(\!\left\lVert f(x_{i,j})\right\rVert_{2}\!{+}\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\!\left\lVert\bar{\phi}{-}\mathbf{X}\nabla\phi\right\rVert_{2}\!\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\!\left\lVert G(x_{i,j})\right\rVert_{2}$
${+}\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert h(x_{i,j})\right\rVert_{2}(\left\lVert\mathbf{S}\right\rVert{+}\left\lVert\tilde{\mathbf{Q}}\right\rVert_{2}\left\lVert J(x_{i,j}\right\rVert_{2})\!\Bigr)$
$\Bigl(\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}\Bigr)+\hat{f}_{i,j}+\frac{1}{2\underline{\pi}_{1}}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j}$
$+\frac{1}{4\underline{\pi}_{1}}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}+\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}\hat{h}_{i,j}.$
@@
Next, we establish conditions for which a quadratic storage function will be found in a ball about the origin for a locally dissipative system.
@@
Finally, we establish when a CPA and quadratic storage function can be determined for a locally dissipative system.
Theorem 17.
Consider Problem 4.2. Let $\mathcal{G}_{\Omega}$ be locally QSR dissipative on $\Omega$ with $\mathbf{Q}\preceq\frac{1}{2}\min(-(\Pi_{2}+\Pi_{3})^{-1},-(\Pi_{7}+\Pi_{8}+2\Pi_{9})^{-1}$ and $\hat{\mathbf{R}}\preceq\frac{1}{2}\min(-(\Pi_{1}^{-1}+\eta\mathbf{I}),-(2\Pi_{4}^{-1}+\Pi_{5}^{-1}+\Pi_{6}^{-1})$, where $\eta\geq\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\hat{J}_{i,j}^{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}$ and $j\in\mathbb{Z}_{0}^{n}.$ Let there exist a storage function $\phi\in\mathcal{C}^{2}$ for which (5) holds in $\Omega,$ where $\sup_{x\in\Omega}\left\lVert\nabla\phi\right\rVert_{2}\leq C$ and $\max_{i,j\in\mathbb{Z}_{1}^{n},x\in\Omega}\left\lvert\frac{\partial\phi}{\partial x^{(i)}\partial x^{(j)}}\Big|_{x}\right\rvert\leq A$ for some $C,A<\infty,$ and let there exist a Lyapunov function $\Phi(x)=x^{\top}\mathbf{P}x$ $(\mathbf{P}\succ 0)$ such that (5) holds for (1) when linearized about the origin and $\Phi(x)\geq\phi(x)$ for all $x\in B_{\epsilon}(0)$. Then, Problem 4.2 will find a valid storage function, $V$ over $\Omega$ for (1) using a non-zero $\epsilon$ and a finite number of simplices.
{pf}
On $B_{\epsilon}(0)$ the theorem holds due to Theorem 16. Next, consider $\Omega$ as a whole. Like step 3 of [giesl2012construction, Theorem 3.3], we construct a continuous storage function by considering the convex combination of two storage functions. Define $\rho:B_{\epsilon}\setminus\Psi\rightarrow\mathbb{R}_{0}^{1}$ as a non-decreasing, continuous function where $\rho(x)=0$ for all $x\in\delta\Psi$ and $\rho(x)=1$ for all $x\in\delta(\Omega\setminus B_{\epsilon}(0)).$ Define the function
$V_{\phi}(x)=\begin{cases}\Phi(x)&x\in\Psi\\ \rho(x)\Phi(x)+(1-\rho(x))\phi(x)&x\in B_{\epsilon}(0)\setminus\Psi\\ \phi(x)&x\in\Omega\setminus B_{\epsilon}(0).\end{cases}$
By assumption, both $\Phi$ and $\phi$ are storage functions. The function $V_{\phi}$ is continuous by the composition of continuous functions. Further, $V_{\phi}$ a storage function as by [willems1972dissipative, Theorem 3] the convex combination of storage functions is also a storage function. Note also that $\sup_{x\in\Omega}\left\lVert\nabla V_{\phi}\right\rVert_{2}\leq C_{\phi}$ and $\max_{i,j\in\mathbb{Z}_{1}^{n},x\in\Omega}\left\lvert\frac{\partial\phi}{\partial x^{(i)}\partial x^{(j)}}\Big|_{x}\right\rvert\leq A_{\phi}$ for some $C_{\phi},A_{\phi}<\infty.$ Thus, by Theorem 15, a viable CPA storage function will be found across $\Omega\setminus B_{\epsilon}(0)$ that adheres to $V_{\phi}(x).$ This function will also adhere to (20a) and (20b) by definition of $V_{\phi}.$ Thus, Problem 4.2 will is able to determine a storage function with finite simplices and a non-zero $\epsilon$ about the origin. ∎